# DIAS: A Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

### v3

@@ IV Methods
Automatic IA segmentation in DSA sequences is an under-explored and challenging task. To fully explore potential research topics, an elaborately designed benchmark for IA segmentation was proposed, aiming to provide comprehensive evaluation of the DIAS dataset using FSS, WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
@@ IV-A Fully-supervised Segmentation
Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network (VSS-Net) efficiently captures spatio-temporal representations of vessels across various scales to achieve precise IA segmentation in DSA. Figure 5 (a,b) illustrates an overview of the proposed network architecture. VSS-Net comprises multi-scale sequence feature extraction modules (SFEM). These modules handle the 2D+Time input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations. Subsequently, these feature maps can be fed into any 2D model following the encoder-decoder framework, enabling vessel segmentation in the sequence.
In the encoder, the input feature map from each stage is integrated with the raw sequence, serving to supplement fine-grained semantic information, except for the initial stage. Specifically, the raw sequence has been resized to be consistent with the corresponding feature maps by bilinear interpolation. Each image of sequence are concatenated with feature maps from the same temporal sequence, and then they are fused using a 1$\times$1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels. This process is formulated as:
$Q=\sum_{i=1}^{N}W\ast(x_{i}\oplus f_{i}),$ (1)
where $x$ represents the input frame, $f$ is the feature map, $N$ denotes the length of the sequence. $W$ are $1\times 1$ convolutional kernels. $\ast$ denotes the convolution operation. $\oplus$ denotes the concatenation operation.
@@
where $\sigma$ represents the sigmoid function. $\tanh$ represents the hyperbolic tangent function. $\odot$ denotes the Hadamard product, i.e., element-wise multiplication. $W_{r}$, $W_{z}$, $W_{h}$ and $U_{r}$, $U_{z}$, $U_{h}$ are 2D-convolutional kernels.
@@ IV-B Weakly-supervised Segmentation
A prevalent strategy for scribble-supervised semantic segmentation is to handle the problem as a fully supervised task, using a loss function $\mathcal{L}$ on the available labeled pixels. However, this baseline approach fails to harness the unlabeled points that contain critical semantic information. Consequently, the absence of sufficient labeled data invariably results in performance degradation. To enhance the segmentation performance of scribble, the key lies in extending the supervision of labeled pixels to unlabeled pixels. To overcome these issues, we propose SSCR, a novel framework for learning-based scribble image segmentation composed of scribble supervision and consistency regularization (CR). Fig. 5 (c) presents the implementation framework for SSCR. SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing segmentation network using mutual consistency and cross pseudo supervision (CPS).
@@
$\mathcal{L}_{scr}=$ $\frac{1}{2}(\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL)),$ (11)
$\mathcal{L}_{cps}=$ $\frac{1}{2}(\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A})),$ (12)
@@
where $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function. $P_{A}$ and $P_{B}$ are the outputs of two models. $PL_{A}$ and $PL_{B}$ are the pseudo labels. $SL$ is the scribble label.
@@ V-A Implementation Details and Evaluation Metrics
TABLE II: Results of fully-supervised segmentation with state-of-the-art models.
Methods Params / M Metrics
[…]
2D UNet [27] 31.04 0.7677 0.9647 0.7570 0.9829 0.6258 0.9811 68.29
Res-UNet [51] 9.37 0.7656 0.9655 0.7425 0.9850 0.6235 0.9816 60.56
UNet++ [52] 9.05 0.7768 0.9659 0.7726 0.9830 0.6374 0.9830 54.65
Att-UNet [53] 8.73 0.7760 0.9660 0.7686 0.9835 0.6258 0.9811 59.37
[…]
MAA-Net [13] 5.80 0.7693 0.9661 0.7423 0.9855 0.6361 0.9818 68.18
FR-UNet [12] 7.37 0.7779 0.9665 0.7646 0.9841 0.6376 0.9827 53.92
3D UNet [27] 90.27 0.7691 0.9648 0.7690 0.9824 0.6275 0.9819 62.92
Res-UNet [51] 26.38 0.7701 0.9653 0.7597 0.9835 0.6284 0.9817 73.05
UNet++ [52] 26.17 0.7649 0.9639 0.7711 0.9811 0.6214 0.9811 67.02
Att-UNet [53] 23.6 0.7677 0.9647 0.7570 0.9829 0.6258 0.9811 68.29
[…]
FR-UNet [12] 20.57 0.7635 0.9646 0.7518 0.9830 0.6210 0.9815 65.67
[…]
PSC [47] 6.43 0.7649 0.9646 0.7552 0.9832 0.6223 0.9814 76.04
SVS-Net [38] 8.48 0.7698 0.9659 0.7474 0.9853 0.6277 0.9803 97.28
ST-UNet [8] 104.37 0.7818 0.9675 0.7533 0.9864 0.6436 0.9843 62.59
VSS-Net 18.69 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
The DIAS dataset was divided into 30 training samples, 10 validation samples, and 20 testing samples on the patient level. To facilitate input into the model, all DSA sequences were resampled to a length of 8. Furthermore, z-score normalization was applied across all DSA sequences for standardization. Patches of size $64\times 64$ were randomly cropped for training, with a batch size of 64. The model parameters were optimized using AdamW [55] with an initial learning rate of 5e-4. The learning rate is gradually reduced by the cosine annealing algorithm over 100 epochs. The parameter $P$ was set to 0.5 in the RPST, i.e., half of the patches in each batch were extracted from labeled data, while the rest came from unlabeled data in SSS experiment. For the SSCR, the weights of loss components $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 1 and 0.5. We only use one network to generate the results for evaluation. All experiments used PyTorch with constant hyperparameters and were conducted on a single GeForce RTX 3090 GPU. We calculated the dice similarity coefficient (DSC), the area under the receiver operating characteristic curve (AUC), accuracy (Acc), sensitivity (Sen), specificity (Spe), and intersection over union (IOU) to evaluate the performance of DSA sequence segmentation. Moreover, we introduced the vascular connectivity (VC) [12] metric, which calculates the ratio of connected components in the segmentation image to those in the ground truth, as a means of measuring the connectivity of vessels.
@@ V-B Fully-supervised Segmentation
where $Squ$ is dimensionality squeeze operation and $\mathcal{H}_{2d}$ is a 2D model.
For a 3D model $\mathcal{H}_{3d}$, the DSA sequence is fed directly into the model, where it is treated as volume data. The dimension, $C$, of the model output is squeezed by DR operation. Finally, $1\times 1$ convolution reduces $F$ to the number of classes:
@@
Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC and VC, among all the evaluated methods. Using FR-UNet as a baseline, VSS-Net shows improvements in numerous metrics except for Spe. Specifically, Sen has increased by 1.3%. ST U-Net, also employing ConvGRU, shows performance very close to VSS-Net, achieving the best results in Acc and Spe. However, the number of model parameters is approximately 5.58 times that of VSS-Net. The VC score that is 19.57 higher than VSS-Net indicates poorer vascular connectivity. Both IPN and SVS-Net employ 3D convolutions in the encoder and 2D convolutions in the decoder, incorporating dimensionality reduction in the intermediate stages. PSC’s segmentation results are similar to those of 3D models, but IPN shows poorer performance, possibly due to the absence of skip connections, which are significantly important for enhancing the detailing in vascular structures.
@@
pCE [47] 0.7455 0.9586 0.7845 0.9735 0.5966 0.9379 60.11 0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 10.02
EM [56] 0.7458 0.9575 0.8029 0.9707 0.5970 0.9725 41.18 0.6356 0.9226 0.8508 0.9281 0.4699 0.9618 9.05
RLoss [45] 0.7146 0.9600 0.6539 0.9867 0.5581 0.9367 46.30 0.5916 0.9147 0.7812 0.9257 0.4232 0.9475 5.17
IVM [46] 0.6606 0.9299 0.8546 0.9360 0.4971 0.9455 71.58 0.6322 0.9198 0.8760 0.9232 0.4662 0.9628 14.78
DMPLS [40] 0.7471 0.9581 0.7963 0.9717 0.5995 0.9708 37.56 0.6327 0.9208 0.8508 0.9259 0.4668 0.9643 9.86
SSCR 0.7589 0.9603 0.8079 0.9732 0.6141 0.9761 36.52 0.6722 0.9343 0.8348 0.9421 0.5115 0.9658 10.30
@@ V-C Weakly-supervised Segmentation
We compared our method with five scribble-supervised segmentation methods: pCE [57], Entropy Minimization (EM) [56], Regularized Loss (RLoss) [45], Intra-class Intensity Variance Minimization (IVM) [46] and Dynamically Mixed Pseudo Labels Supervision (DMPLS) [40]. To ensure a fair comparison, all experimental networks employed the UNet [27], consistent with the approach in DMPLS. Table III presents the results of scribble-supervised segmentation of the DIAS dataset. Fig. 7 showcases the visual results of all methods. Thanks to pseudo supervision and consistency regularization for scribble supervision, SSCR achieved significantly better performance in RDFA and SALE by obtaining the highest scores in most metrics among all methods. It is noteworthy that in the segmentation results of IVM, the Sen on RDFA is 4.67% higher compared to SSCR, and it is 4.12% higher on SALE. Sen is used to measure the proportion of correctly identified positive classes (i.e., vessels) out of all actual positive classes, independent of the background. The visualization results from IVM indicate that achieving high Sen comes at the cost of introducing numerous incorrectly segmented vessel pixels. Consequently, this leads to lower scores in other evaluation metrics such as the DSC and AUC. Regarding SALE, the vessels segmented by the same method are thicker than the ones trained with RDFA annotation, leading to a larger number of false positive pixels at the vessel’s edge. This outcome may have been caused by the SALE focusing on the annotation of thick vessels. Similarly, due to the lack of supervision for small vessels, where vascular connectivity issues are more prevalent, SALE struggles to extract fine vessels, resulting in a VC score significantly better than RDFA. From the last column of the figure, it is evident that SSCR’s segmentation results have the least number of blue pixels, corroborating the best quantitative results presented in the Table III.
@@
0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 10.02
0.6364 0.9244 0.8346 0.9316 0.4706 0.9479 6.57
0.6472 0.9275 0.8385 0.9346 0.4827 0.9445 7.29
0.6484 0.9276 0.8536 0.9334 0.4827 0.9639 8.28
0.6722 0.9343 0.8348 0.9421 0.5115 0.9658 10.30
TABLE V: Ablation research results of RPST for semi-supervised IA sequence segmentation.
@@
Teacher 1 0 0.6266 0.9259 0.6853 0.9451 0.4693 0.8538 385.62
Student 1 30 0.6486 0.9294 0.7085 0.9464 0.4941 0.9505 107.35
S w/o SDA 1 60 0.6394 0.9296 0.6922 0.9483 0.4832 0.9352 108.96
Student 1 60 0.6502 0.9305 0.7177 0.9482 0.4958 0.9433 103.09
Teacher 3 0 0.6919 0.9490 0.7358 0.9673 0.5336 0.9182 320.02
Student 3 30 0.7265 0.9555 0.7478 0.9721 0.5745 0.9591 121.84
S w/o SDA 3 60 0.7228 0.9555 0.7456 0.9734 0.5700 0.9578 125.83
Student 3 60 0.7275 0.9549 0.7595 0.9778 0.5893 0.9616 125.78
Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.5910 0.9694 148.89
Student 10 30 0.7544 0.9618 0.7660 0.9786 0.6087 0.9769 93.31
S w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 93.92
Student 10 60 0.7579 0.9625 0.7662 0.9795 0.6130 0.9786 87.72
We conducted ablation studies to investigate the impact of various components in our proposed framework for weakly-supervised IA segmentation. As illustrated in Table IV, with pCE serving as the baseline, we incrementally added components for comparative analysis. Initially, the introduction of pseudo-label supervision resulted in a slight decline in metrics such as DSC and AUC. Based on this, we utilized a dual-model framework to establish cross pseudo-supervision, leading to a 1.08% enhancement in DSC. Compared to a single-model approach, a dual-model setup, receiving the same scribble supervision, generates pseudo-labels for mutual supervision. This avoids the scenario of a model being self-supervised by its own predictions, thereby enhancing the model’s generalizability. We employ consistency regularization on the pseudo-labels generated by dual models to penalize inconsistent segmentation. Compared to ‘pCE+PS’, ‘pCE+PS+CR’ exhibits a 1.2% improvement in DSC. Additionally, it has attained the highest Sen. We implement a combined strategy of cross pseudo-supervision and consistency regularization on the pseudo-labels produced by dual models. The integrated approach, termed as SSCR and denoted as ‘pCE+PS+CPS+CR’, has achieved notable improvements, surpassing ‘pCE+PS+CPS’ and ‘pCE+PS+CR’ by 2.5% and 2.38%, respectively.

### v4

@@ IV Methods
Automatic IA segmentation in DSA sequences is an under-explored and challenging task. To fully explore potential research topics, an elaborately designed benchmark for IA segmentation was proposed, aiming to provide comprehensive evaluation of the DIAS dataset using Fully-Supervised Segmentation (FSS), WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
@@ IV-A Fully-supervised Segmentation
Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network efficiently captures spatio-temporal representations of vessels across various scales to achieve precise IA segmentation in DSA. Figure 5 (a,b) illustrates an overview of the proposed network architecture. In VSS-Net, we employ Sequence Feature Extraction Modules (SFEM) at each stage in the encoder to extract feature maps across multiple scales. These modules handle the 2D+Time input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations. With the deepening of stages, the input resolution is reduced, consequently expanding the receptive field that facilitates the generation of high-dimensional spatiotemporal representations. Subsequently, these feature maps can be fed into any 2D model following the encoder-decoder framework, enabling vessel segmentation in the sequence.
In the encoder, the input feature map from each stage is integrated with the raw sequence, serving to supplement fine-grained semantic information, except for the initial stage. Specifically, the raw sequence has been resized to be consistent with the corresponding feature maps by bilinear interpolation. Each image in the sequence is concatenated with feature maps from the same temporal sequence, and then they are fused using a 1$\times$1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels. The calculation process for the fusion output $Q^{j}$ at scale $j$ is as follows:
$Q^{j}=\sum_{i=1}^{N}W\ast(x^{j}_{i}\oplus f^{j}_{i}),\quad\text{for }j\in\left\{\frac{1}{2},\frac{1}{4},\ldots,\frac{1}{2^{n-1}}\right\},$ (1)
where $x^{j}_{i}$ represents the $i^{th}$ frame of the input sequence, scaled to $j$ times the original resolution, $f$ is the feature map, $N$ denotes the length of the sequence, $W$ are $1\times 1$ convolutional kernels, $\ast$ denotes the convolution operation, $\oplus$ denotes the concatenation operation, $n$ denotes the number of stages in the encoder.
@@
where $\sigma$ represents the sigmoid function, $\tanh$ represents the hyperbolic tangent function, $\odot$ denotes the Hadamard product, i.e., element-wise multiplication, $W_{r}$, $W_{z}$, $W_{h}$ and $U_{r}$, $U_{z}$, $U_{h}$ are 2D-convolutional kernels.
@@ IV-B Weakly-supervised Segmentation
A prevalent strategy for scribble-supervised semantic segmentation is to handle the problem as a fully supervised task, using a loss function $\mathcal{L}$ on the available labeled pixels. However, this baseline approach fails to harness the unlabeled points that contain critical semantic information. Consequently, the absence of sufficient labeled data invariably results in performance degradation. To enhance the segmentation performance of scribble, the key lies in extending the supervision of labeled pixels to unlabeled pixels. To overcome these issues, we propose SSCR, a novel framework for learning-based scribble image segmentation composed of scribble supervision and consistency regularization. Fig. 5 (c) presents the implementation framework for SSCR. SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing the segmentation network using mutual consistency and cross pseudo supervision.
@@
$\mathcal{L}_{scr}=$ $\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL),$ (11)
$\mathcal{L}_{cps}=$ $\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A}),$ (12)
@@
where $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function, $P_{A}$ and $P_{B}$ are the outputs of two models, $PL_{A}$ and $PL_{B}$ are the pseudo labels, $SL$ is the scribble label.
@@ V-A Implementation Details and Evaluation Metrics
The DIAS dataset was divided into 30 training samples, 10 validation samples, and 20 testing samples on the patient level. To facilitate input into the model, all DSA sequences were resampled to a length of 8. Furthermore, z-score normalization was applied across all DSA sequences for standardization. Patches of size $64\times 64$ were randomly cropped for training, with a batch size of 64. The model parameters were optimized using AdamW [57] with an initial learning rate of 5e-4. The learning rate is gradually reduced by the cosine annealing algorithm over 100 epochs. The parameter $P$ was set to 0.5 in the RPST, i.e., half of the patches in each batch were extracted from labeled data, while the rest came from unlabeled data in SSS experiment. For the SSCR, the weights of loss components $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 1 and 0.5. We only use one network to generate the results for evaluation. All experiments used PyTorch with constant hyperparameters and were conducted on a single GeForce RTX 3090 GPU. We calculated the Dice Similarity Coefficient (DSC), the Area Under the receiver operating Characteristic curve (AUC), accuracy (Acc), sensitivity (Sen), specificity (Spe), and Intersection Over Union (IOU) to evaluate the performance of DSA sequence segmentation. Furthermore, we conducted evaluations of the Vascular Connectivity (VC) metric [13] and the connectivity-aware similarity measure (clDice) [58].
@@ V-B Fully-supervised Segmentation
where $Squ$ is dimensionality squeeze operation and $\mathcal{H}_{2d}$ is a 2D model. For a 3D model $\mathcal{H}_{3d}$, the DSA sequence is fed directly into the model, where it is treated as volume data. The dimension, $C$, of the model output is squeezed by DR operation. Finally, $1\times 1$ convolution reduces $F$ to the number of classes:
@@
[…]
Methods Params / M Metrics
[…]
2D UNet 31.04 0.7700$\pm$0.0525 0.9654$\pm$0.0125 0.7570$\pm$0.0879 0.9839$\pm$0.0098 0.6290$\pm$0.0695 0.9829$\pm$0.0080 0.6939$\pm$0.0628 63.60$\pm$39.01
Res-UNet 9.37 0.7656$\pm$0.0548 0.9655$\pm$0.0120 0.7425$\pm$0.0950 0.9850$\pm$0.0083 0.6235$\pm$0.0719 0.9816$\pm$0.0093 0.6870$\pm$0.0745 60.51$\pm$40.00
UNet++ 9.05 0.7767$\pm$0.0468 0.9659$\pm$0.0122 0.7726$\pm$0.0752 0.9830$\pm$0.0091 0.6374$\pm$0.0630 0.9830$\pm$0.0084 0.7042$\pm$0.0589 54.79$\pm$32.24
Att-UNet 8.73 0.7760$\pm$0.0435 0.9660$\pm$0.0118 0.7686$\pm$0.0800 0.9835$\pm$0.0092 0.6361$\pm$0.0587 0.9837$\pm$0.0077 0.7030$\pm$0.0569 59.39$\pm$31.50
CS2-Net 8.93 0.7747$\pm$0.0436 0.9664$\pm$0.0120 0.7524$\pm$0.0878 0.9855$\pm$0.0078 0.6344$\pm$0.0586 0.9833$\pm$0.0083 0.6941$\pm$0.0588 64.91$\pm$45.90
MAA-Net 5.80 0.7693$\pm$0.0508 0.9661$\pm$0.0120 0.7423$\pm$0.0969 0.9855$\pm$0.0091 0.6279$\pm$0.0669 0.9818$\pm$0.0095 0.69$\pm$0.0690 68.20$\pm$42.60
FR-UNet 7.37 0.7779$\pm$0.0459 0.9665$\pm$0.0118 0.7646$\pm$0.0808 0.9841$\pm$0.0083 0.6376$\pm$0.0615 0.9827$\pm$0.0078 0.7112$\pm$0.0603 53.92$\pm$27.93
3D UNet 90.27 0.7692$\pm$0.0490 0.9648$\pm$0.0121 0.7690$\pm$0.0887 0.9824$\pm$0.0090 0.6275$\pm$0.0650 0.9818$\pm$0.0090 0.6975$\pm$0.0663 62.95$\pm$37.21
Res-UNet 26.38 0.7701$\pm$0.0454 0.9653$\pm$0.0121 0.7597$\pm$0.0809 0.9835$\pm$0.0087 0.6284$\pm$0.0606 0.9817$\pm$0.0092 0.6991$\pm$0.0606 73.19$\pm$45.50
UNet++ 26.17 0.7649$\pm$0.0448 0.9639$\pm$0.0121 0.7712$\pm$0.0791 0.9811$\pm$0.0088 0.6215$\pm$0.0589 0.9811$\pm$0.0092 0.6989$\pm$0.0595 67.03$\pm$48.51
Att-UNet 23.60 0.7677$\pm$0.0513 0.9647$\pm$0.0131 0.7570$\pm$0.0816 0.9829$\pm$0.0106 0.6258$\pm$0.0686 0.9811$\pm$0.0098 0.6966$\pm$0.0624 68.32$\pm$40.96
CS2-Net 23.83 0.7687$\pm$0.0472 0.9645$\pm$0.0124 0.7712$\pm$0.0838 0.9816$\pm$0.0099 0.6267$\pm$0.0628 0.9808$\pm$0.0100 0.6960$\pm$0.0650 67.91$\pm$45.25
FR-UNet 20.57 0.7635$\pm$0.0587 0.9646$\pm$0.0123 0.7518$\pm$0.0889 0.9830$\pm$0.0107 0.6211$\pm$0.0754 0.9815$\pm$0.0096 0.6913$\pm$0.0686 65.65$\pm$45.09
[…]
PSC 6.43 0.7649$\pm$0.0523 0.9646$\pm$0.0123 0.7552$\pm$0.0896 0.9832$\pm$0.0095 0.6223$\pm$0.069 0.9814$\pm$0.0104 0.6965$\pm$0.0604 76.05$\pm$44.63
SVS-Net 8.48 0.7697$\pm$0.0435 0.9659$\pm$0.0115 0.7474$\pm$0.0826 0.9853$\pm$0.0072 0.6277$\pm$0.0578 0.9803$\pm$0.0085 0.6880$\pm$0.0580 97.33$\pm$49.10
ST-UNet 104.37 0.7818$\pm$0.0406 0.9675$\pm$0.0119 0.7533$\pm$0.0821 0.9864$\pm$0.0070 0.6435$\pm$0.0551 0.9843$\pm$0.0075 0.7073$\pm$0.0587 62.54$\pm$38.25
VSS-Net 18.69 0.7822$\pm$0.0419 0.9668 $\pm$ 0.0119 0.7770$\pm$0.0830 0.9838$\pm$0.0076 0.6442$\pm$0.0568 0.9857$\pm$0.0070 0.7119$\pm$0.0583 43.02$\pm$27.47
Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC, clDice and VC, among all the evaluated methods. Using 2D FR-UNet as a baseline, VSS-Net shows improvements in numerous metrics except for Spe. Specifically, Sen has increased by 1.24%. ST U-Net, also employing ConvGRU, shows performance very close to VSS-Net, achieving the best results in Acc and Spe. However, the number of model parameters is approximately 5.58 times that of VSS-Net. The VC score that is 19.52 higher than VSS-Net indicates poorer vascular connectivity. Both IPN and SVS-Net employ 3D convolutions in the encoder and 2D convolutions in the decoder, incorporating dimensionality reduction in the intermediate stages. PSC’s segmentation results are similar to those of 3D models, but IPN shows poorer performance, possibly due to the absence of skip connections, which are significantly important for enhancing the detailing in vascular structures.
Figure 6: Dimensionality reduction operation applied to 2D/3D models
@@
pCE 0.7455 0.9586 0.7845 0.9735 0.5966 0.9379 0.7022 60.11 0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 0.6589 10.02
EM 0.7458 0.9575 0.8029 0.9707 0.5970 0.9725 0.7103 41.18 0.6356 0.9226 0.8508 0.9281 0.4699 0.9618 0.6576 9.05
RLoss 0.7146 0.9600 0.6539 0.9867 0.5581 0.9367 0.5906 46.30 0.5916 0.9147 0.7812 0.9257 0.4232 0.9475 0.5702 5.17
IVM 0.6606 0.9299 0.8546 0.9360 0.4971 0.9455 0.6910 71.58 0.6322 0.9198 0.8760 0.9232 0.4662 0.9628 0.6738 14.78
DMPLS 0.7471 0.9581 0.7963 0.9717 0.5995 0.9708 0.6996 37.56 0.6327 0.9208 0.8508 0.9259 0.4668 0.9643 0.6601 9.86
SSCR 0.7651 0.9639 0.7620 0.9810 0.6214 0.9778 0.6977 34.48 0.6574 0.9277 0.8398 0.9344 0.4958 0.9672 0.6732 6.47
@@ V-C Weakly-supervised Segmentation
We compared our method with five scribble-supervised segmentation methods: pCE [63], Entropy Minimization (EM) [64], Regularized Loss (RLoss) [51], Intra-class Intensity Variance Minimization (IVM) [52] and Dynamically Mixed Pseudo Labels Supervision (DMPLS) [41]. To ensure a fair comparison, all experimental networks employed the UNet [28], consistent with the approach in DMPLS. Table III presents the results of scribble-supervised segmentation of the DIAS dataset. Fig. 8 showcases the visual results of all methods. Thanks to pseudo supervision and consistency regularization for scribble supervision, SSCR achieved significantly better performance in RDFA and SALE by obtaining the highest scores in most metrics among all methods. It is noteworthy that in the segmentation results of IVM, the Sen on RDFA is 9.26% higher compared to SSCR, and it is 3.62% higher on SALE. Sen is used to measure the proportion of correctly identified positive classes (i.e., vessels) out of all actual positive classes, independent of the background. The visualization results from IVM indicate that achieving high Sen comes at the cost of introducing numerous incorrectly segmented vessel pixels. Consequently, this leads to lower scores in other evaluation metrics such as the DSC and AUC. Regarding SALE, the vessels segmented by the same method are thicker than the ones trained with RDFA annotation, leading to a larger number of false positive pixels at the vessel’s edge. This outcome may have been caused by the SALE focusing on the annotation of thick vessels. Similarly, due to the lack of supervision for small vessels, where vascular connectivity issues are more prevalent, SALE struggles to extract fine vessels, resulting in a VC score significantly better than RDFA. From the last column of the figure, it is evident that SSCR’s segmentation results have the least number of blue pixels, corroborating the best quantitative results presented in the Table III.
@@
0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 0.6589 10.02
0.6364 0.9244 0.8346 0.9316 0.4706 0.9479 0.6437 6.57
0.6472 0.9275 0.8385 0.9346 0.4827 0.9445 0.6692 7.29
0.6484 0.9276 0.8536 0.9334 0.4827 0.9639 0.6685 8.28
0.6574 0.9277 0.8398 0.9344 0.4958 0.9672 0.6732 6.47
TABLE V: Ablation research results of RPST for semi-supervised IA sequence segmentation. (L: Labeled data, U: Unlabeled data)
@@
Teacher 1 0 0.6266 0.9259 0.6853 0.9451 0.4693 0.8538 0.5173 385.62
Student 1 30 0.6486 0.9294 0.7085 0.9464 0.4941 0.9505 0.5668 107.35
S w/o SDA 1 60 0.6394 0.9296 0.6922 0.9483 0.4832 0.9352 0.5531 108.96
Student 1 60 0.6502 0.9305 0.7177 0.9482 0.4958 0.9433 0.5653 103.09
Teacher 3 0 0.6919 0.9490 0.7358 0.9673 0.5336 0.9182 0.5921 320.02
Student 3 30 0.7265 0.9555 0.7478 0.9721 0.5745 0.9591 0.6523 121.84
S w/o SDA 3 60 0.7228 0.9555 0.7456 0.9734 0.5700 0.9578 0.6404 125.83
Student 3 60 0.7275 0.9549 0.7595 0.9778 0.5893 0.9616 0.6502 125.78
Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.5910 0.9694 0.6512 148.89
Student 10 30 0.7544 0.9618 0.7660 0.9786 0.6087 0.9769 0.6747 93.31
S w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 0.6708 93.92
Student 10 60 0.7579 0.9625 0.7662 0.9795 0.6130 0.9786 0.6794 87.72
We conducted ablation studies to investigate the impact of various components in our proposed framework for weakly-supervised IA segmentation. As illustrated in Table IV, with pCE serving as the baseline, we incrementally added components for comparative analysis. Initially, the introduction of pseudo-label supervision resulted in a slight decline in metrics such as Sen, AUC and clDice. Based on this, we utilized a dual-model framework to establish cross pseudo-supervision, leading to a 1.08% enhancement in DSC. Compared to a single-model approach, a dual-model setup, receiving the same scribble supervision, generates pseudo-labels for mutual supervision. This avoids the scenario of a model being self-supervised by its own predictions, thereby enhancing the model’s generalizability. We employ consistency regularization on the pseudo-labels generated by dual models to penalize inconsistent segmentation. Compared to ‘pCE+PS’, ‘pCE+PS+CR’ exhibits a 1.2% improvement in DSC. Additionally, it has attained the highest Sen. We implement a combined strategy of cross pseudo-supervision and consistency regularization on the pseudo-labels produced by dual models. The integrated approach, termed as SSCR and denoted as ‘pCE+PS+CPS+CR’, has achieved notable improvements, surpassing ‘pCE+PS+CPS’ and ‘pCE+PS+CR’ by 1.02% and 0.9%, respectively.